4. assume that GDp is $6,000, personal disposable income is $5,100, and the government budget deficit is $200. Consumption is $3,800, and the trade deficit is $100. a. how large is saving (S)? b. how large is investment (I)? c. how large is government spending (G)?


1
Expert's answer
2021-05-11T09:05:39-0400

(a) Saving equals to the government savings plus private savings:


"S=GS+PS."

Private savings equals to disposable income minus consumption:


"PS=DI-C=\$5100-\$3800=\$1300."

The GS equals -$200 (the government spent more than it earned) so we can find saving:


"S=GS+PS=-\$200+\$1300=\$1100."

(b)-(c) Let's first find the government spending (G). From the condition of the question we know that the government budget deficit is $200:


"T-G=-\$200."

We can find taxes (T) as follows:


"T=GDP-DI=\$6000-\$5100=\$900."

Then, we can find government spending:


"G=T+\$200=\$900+\$200=\$1100."

Finally, we can find the  investment as follows:


"I=Y-C-G-NX,""I=\$6000-\$3800-\$1100-(-\$100)=\$1200."